Directions: identify the test tool use in a certain problem



1. Hospital Infections A medical investigation claims that the average number of infections per week at a hospital in south-western Pennsylvania is 16.3. A random sample of 10 weeks had a mean number of 17.7 infections. The sample standard deviation is 1.8. Is there enough evidence to reject the investigator's claim at alpha = 0.05 ?




Type of test:




Reason/s:

1
Expert's answer
2022-04-26T10:44:24-0400

This is a one-sample t-test.

"H_0: \mu=16.3"

"H_1: \mu \not=16.3"

Test statistic: "t=\frac{17.7-16.3}{1.8/\sqrt{10}}=2.46"

P-value: p=2P(T>2.46)=0.0362.


Since the p-value is less than 0.05, reject the null hypothesis.


There is enough evidence to reject the investigator’s claim at α = 0.05.